# Plasticity and the corresponding mechanisms of nanopowdered Mg during shock consolidation and spallation

### 01 Mar 2024 v1

@@ III.2 Release and tension stage
When the shock wave travels to the free surface, it then triggers a release wave due to the shock reflection (e.g., release stage), propagating along the opposite direction (right left), accelerating an apparent decrement of stress (R, Fig. 5). The subsequent interaction between two release waves, induces a tension wave within the material, causing a negative stress, i.e., the tensile stress (e.g., tension stage, T, Fig. 5). When the tensile stress exceeds the ultimate strength of the material, it facilitates an apparent spallation.(Huang et al., 2012) However, there still exists two most significant questions in the shock dynamics: i) Is the deformation (i.e., deformation twinning) reversible when the release wave passes? What is the mechanism for the spallation caused by stretching? It is difficult to acquire the answers accurately, via analyzing the postmortem microstructures in the traditional dynamic experiments, owing to its limited information on the process.(Dixit et al., 2017; Kannan et al., 2018)

### 13 Mar 2024 v2

@@ III.2 Release and tension stage
When the shock wave travels to the free surface, it then triggers a release wave due to the shock reflection (e.g., release stage), propagating along the opposite direction (right left), accelerating an apparent decrement of stress (R, Fig. 5). The subsequent interaction between two release waves, induces a tension wave within the material, causing a negative stress, i.e., the tensile stress (e.g., tension stage, T, Fig. 5). When the tensile stress exceeds the ultimate strength of the material, it facilitates an apparent spallation.Huang et al. (2012) However, there still exists two most significant questions in the shock dynamics: i) Is the deformation (i.e., deformation twinning) reversible when the release wave passes? What is the mechanism for the spallation caused by stretching? It is difficult to acquire the answers accurately, via analyzing the postmortem microstructures in the traditional dynamic experiments, owing to its limited information on the process.Dixit et al. (2017); Kannan et al. (2018)